# A Diffusion-Based Approach for Simulating Forward-in-Time State-Dependent Speciation and Extinction Dynamics

### v1

@@ 1 Introduction
The branching events of a phylogenetic tree exhibit a pattern that stores information about the underlying speciation and extinction processes [29]. Nee et al. [29] treated both speciation and extinction as a constant-rate birth-death process by which lineages give birth to new lineages (speciation) at a rate $\lambda$ and lineages die (extinction) at a rate $\mu$. Speciation and extinction rates, however, are expected to vary idiosyncratically among phylogenetic lineages and over geological timescales. Workers have designed birth-death models to study a variety of intrinsic and extrinsic factors that might shape diversification rates. Species age [1, 11, 32] and inherited traits [6, 7, 18, 25, 31] are two types of intrinsic factors thought to drive diversification rates, whereas environment  [4, 30] and geography [10, 21, 34] are common extrinsic factors of interest. In the end, a common goal of these models is to infer the underlying event rates given an observed phylogenetic pattern either through likelihood-based [27, 33, 24] or likelihood-free approaches [13, 20, 29, 35, 39].
Fundamentally, birth-death processes model the random arrival times of discrete events that generate or “build” a phylogenetic tree over time [29, 25]. As an alternative to this tree-based representation of the process, recent work [3] introduced an equivalent diffusion-based representation for a class of birth-death models with state-dependent rates, known as state-dependent speciation-extinction (SSE) models [25]. As noted by [3], population genetics theory has benefited immensely from diffusion-based approximations to population-based models of allele frequency change, diffusion-based approximations of birth-death models remain underexplored in the phylogenetics literature. Despite the widespread popularity of birth-death models among evolutionary biologists, these models recently entered a phase of intense but overdue scrutiny to better understand what the models can and cannot estimate reliably when fitted to real biological datasets [5, 37, 23, 19, 22, 28]. This has created demand for new frameworks to understand the mathematical properties of these complex stochastic processes to guide biological research programs.
@@ 2.1 Overview of state-dependent speciation and extinction models
In general, SSE models are stochastic branching processes with state-dependent birth (speciation) and death (extinction) rates. The states can either be discrete or continuous [25, 6, 7] and can represent various things, ranging from phenotypic traits to geographical ranges [10]. Some SSE models have processes that are only defined by anagenetic process and state-dependent diversification process [25], while others have processes that are defined by both anagenetic and cladogenetic processes [10, 9] shown in Fig. 1. An anagenetic process is defined as a process of trait evolution within lineages, between branching events. In the BiSSE model [25], this corresponds to trait transition events of going from a discrete trait $A$ to another discrete trait $B$ or vice versa. These trait-dependent transition rates are encoded in the infinitesimal rate matrix $Q$, for which the off-diagonal entry $q_{ij}$ defines the rate of transitioning from state $i$ to $j$. A cladogenetic process is defined as a process in which state transition occurs in conjunction with a branching event (with speciation) of a lineage. SSE models with anagenetic and cladogenetic events are referred to as ClaSSE models.
Figure 1: From left to right: a speciation event without cladogenetic state changes, a speciation event with cladogenetic state changes, an anagenetic state change.
@@
Figure 2: An illustration of GeoSSE events on a phylogeny with range state information.
@@ 2.2 Transforming a stochastic process
The infinitesimal mean and variance parameters for $\{X(t)\}$ are given by,
@@ 2.3 Diffusion-based framework for state-dependent diversification model
In this section, we establish the framework for simulating state dynamics for state-dependent speciation and extinction models using diffusion processes. We show how to implement the framework in the ClaSSE model introduced in [9]. Then, we relate our framework to earlier research [3] using a diffusion process for the BiSSE model [25]. Furthermore, in Section 2.4, we apply our framework to the GeoSSE model [10] and derive results in the later sections.
@@
$\mathbb{P}_{i}^{+}$ $=$ $S_{i}^{+}+E_{i}^{+}+Q_{i}^{+},$
$\mathbb{P}_{i}^{-}$ $=$ $S_{i}^{-}+E_{i}^{-}+Q_{i}^{-},$
$\mathbb{P}_{i}$ $=$ $1-\left[\mathbb{P}_{i}^{+}+\mathbb{P}_{i}^{-}\right],$
@@
$S_{i}^{+}$ $=$ Probability of events that lead to an increase in the number of species in state $i$ through state-dependent
speciation and speciation in conjunction with cladogenetic state change.
$E_{i}^{+}$ $=$ Probability of events that lead to an increase in the number of species in state $i$ through extinction.
$Q_{i}^{+}$ $=$ Probability of events that lead to an increase in the number of species in state $i$ through anagenetic
state change.
$S_{i}^{-}$ $=$ Probability of events that lead to a decrease in the number of species in state $i$ through state-dependent
speciation and speciation in conjunction with cladogenetic state change.
$E_{i}^{-}$ $=$ Probability of events that lead to a decrease in the number of species in state $i$ through extinction.
$Q_{i}^{-}$ $=$ Probability of events that lead to a decrease in the number of species in state $i$ through anagenetic
state change.
@@
To demonstrate the generality of the framework, we show the BiSSE model [25] (and similarly for the MuSSE model [7]) can be represented as a diffusion process as follows. Under the BiSSE model, species possess binary traits with values in the state space $S\in\{1,2\}$. BiSSE is a special case of the ClaSSE model that, while it allows anagenetic trait transition and extinction events, its speciation events do not cause cladogenetic trait changes. That is, daughter lineages identically inherit the parent lineage state following speciation. Readers can refer to the supplementary material from [9] for its derivation. For the BiSSE model, we have
$S_{1}^{+}=\lambda_{1}N_{1},\>E_{1}^{+}=0,\>Q_{1}^{+}=q_{21}N_{2},$
$S_{1}^{-}=0,\>E_{1}^{-}=\mu_{1}N_{1},\>Q_{1}^{-}=q_{12}N_{1},$
where $\lambda_{1}$ and $\mu_{1}$ are speciation and extinction rates for trait $1$, respectively. $q_{12}$ and $q_{21}$
are anagenetic trait transition from $1$ to $2$ and from $2$ to $1$, respectively.
@@ 2.4 Diffusion-based framework for the GeoSSE model
In this section, we use the framework established in Section 2.3 to the GeoSSE model. We present the method for simulating the state dynamics under the diffusion framework in Section 2.5, and derive theoretical results regarding stationary distribution under our framework for the GeoSSE model in Sections 2.6-2.8. The procedure we apply here is compatible with any model within the ClaSSE family.
Consider an $n$-region GeoSSE model where $n\in\mathbb{Z}^{+}$, we define the following state space and variable,
@@
with ranges i and j, respectively. Note that $\text{with ranges $i$ and $j$, respectively. Note that }b^{i}_{j}\equiv b^{j}_{i}.$
@@
We define a stochastic process $\{N_{i}(t)\}$ with infinitesimal mean $\mu_{i}=\mathbb{E}(dN_{i}/dt)$ and variance $\sigma^{2}_{i}=var(dN_{i}/dt)$. Here, $N_{i}(t)$ represents the number of species with range state $i$ at time $t$. The infinitesimal mean $\mu_{i}$ and variance $\sigma^{2}_{i}$ follow directly from Lemma 2. We derive the transition probabilities described in Eq. LABEL:eq:transition in the context of the GeoSSE model, as shown in Eqs. 26-38.
@@
$\mathbb{P}_{i}^{+}$ $=$ $W_{i}^{+}+D_{i}^{+}+B_{i}^{+}+E_{i}^{+}$ (26)
$=$ $\underbrace{\sum_{\mathclap{\begin{subarray}{c}j\in S\end{subarray}}}\sum_{\mathclap{\begin{subarray}{c}\ell\in j\\ \{\ell\}=i\end{subarray}}}N_{j}w_{\ell}}_{W_{i}^{+}}+\underbrace{\sum_{\mathclap{\begin{subarray}{c}k\in i\end{subarray}}}\sum_{\mathclap{\begin{subarray}{c}\ell\in i\\ \ell\neq k\end{subarray}}}N_{i\backslash\{\ell\}}d_{k\ell}}_{D_{i}^{+}}+\underbrace{\sum_{\mathclap{\begin{subarray}{c}j\in S\\ i\subset j\end{subarray}}}N_{j}b^{i}_{j\backslash i}}_{B_{i}^{+}}+\underbrace{\sum_{\begin{subarray}{c}j\in S\\ |j\backslash i|=1\end{subarray}}\sum_{\begin{subarray}{c}\ell\in j\backslash i\end{subarray}}N_{j}e_{\ell}}_{E_{i}^{+}}$
$\mathbb{P}_{i}^{-}$ $=$ $W_{i}^{-}+D_{i}^{-}+B_{i}^{-}+E_{i}^{-}$ (37)
@@
$\mathbb{P}_{i}$ $=$ $1-\left[\mathbb{P}_{i}^{+}+\mathbb{P}_{i}^{-}\right].$ (38)
For clarity, we provide the biogeographic interpretation on how each term in Eqs. 26-38 is derived
@@
    $\bm{W_{i}^{+}}$. To gain a new species with range state $i$ through a within-region speciation event, the new species range $i$ must contain only region $\ell$ ($\ell\in i$ and $|i|=1$). This endemic species can undergo a speciation event with probability $w_{\ell}N_{i}$. Any species with range state $j$ that also occupies region $\ell$ can undergo a within-region speciation event with probability $w_{\ell}\sum_{j\in S}\mathbbm{1}_{i\subseteq j}N_{j}$. The total sum of probabilities for $\bm{W_{i}^{+}}$ is,
    $\sum_{\mathclap{\begin{subarray}{c}j\in S\end{subarray}}}\sum_{\mathclap{\begin{subarray}{c}\ell\in j\\ \{\ell\}=i\end{subarray}}}N_{j}w_{\ell}.$
@@
    $\bm{D_{i}^{+}}$. To gain a new species with range state $i$ through a dispersal event, the species adds the new region $\ell$ to its ancestral range. Species are always widespread immediately following dispersal. The total probability of this event occurring is,
    $\sum_{\mathclap{\begin{subarray}{c}k\in i\end{subarray}}}\sum_{\mathclap{\begin{subarray}{c}\ell\in i\\ \ell\neq k\end{subarray}}}N_{i\backslash\{\ell\}}d_{k\ell}.$
@@
    $\bm{B_{i}^{+}}$. To gain a new species with range state $i$ through a between-region speciation event, the new species can be either endemic or widespread $|i|>0$ that originated from a widespread ancestral species with larger range state $j$ ($i\subset j$). In general, we have no information of whether the new species occurs in left or right lineage following a speciation event, so we do not consider the orientation. The total probability of this event occurring is,
    $\sum_{\mathclap{\begin{subarray}{c}j\in S\\ i\subset j\end{subarray}}}N_{j}b^{i}_{j\backslash i}.$
@@
    $\bm{E_{i}^{+}}$. To gain a new species with range state $i$ through a local extinction event, the ancestral species must have a larger range state $j$ with size that differs by 1 from the new species’ range state $i$ such that $|j\backslash i|=1$. The total probability of this occurring is,
    $\sum_{\begin{subarray}{c}j\in S\\ |j\backslash i|=1\end{subarray}}\sum_{\begin{subarray}{c}\ell\in j\backslash i\end{subarray}}N_{j}e_{\ell}.$
@@
    $\bm{D_{i}^{-}}$. To lose a species with range state $i$ through a dispersal event, the species must disperse to a new region. The species count remains unchanged if the species already occupies all regions ($|i|=|R|$). The probability of this event occurring is,
@@
    $\bm{B_{i}^{-}}$. To lose a species with range state $i$ through a between-region speciation event, the species must be widespread and undergo a speciation event that gives rise to a new species in state $j$ with smaller range state size ($|j|<|i|$). The factor of 1/2 corrects for double-counting the new species with range $j$ being either the left daughter or right daughter lineage. The probability of this event occurring is,
    $\sum_{\begin{subarray}{c}j\in S\\ |j|<|i|\end{subarray}}\frac{1}{2}N_{i}b^{j}_{i\backslash j}.$
@@
    $\bm{E_{i}^{-}}$. To lose a species with range state $i$ through a local extinction event, a species must undergo an extinction event in one of its regions. If the species is endemic, this event leads to total extinction of the species. The probability of this event occurring is,
    $\sum_{\begin{subarray}{c}\ell\in R\\ \ell\in i\end{subarray}}N_{i}e_{\ell}.$
The next section uses Eqs. 26-38 to define the stochastic process $\{\Pi_{i}(t):t>0\}$ that models the frequency of species in range state $i$ at time t. The infinitesimal mean $\mu_{\Pi_{i}}$ and variance $\sigma^{2}_{\Pi_{i}}$ follow directly from Lemma 3.

### v2

@@ 1 Introduction
The branching events of a phylogenetic tree exhibit a pattern that stores information about the underlying speciation and extinction processes [29]. In [29], they first considered a model where both speciation and extinction are treated as a constant-rate birth-death process by which lineages give birth to new lineages (speciation) at a rate $\lambda$ and lineages die (extinction) at a rate $\mu$. Speciation and extinction rates, however, are expected to vary idiosyncratically among phylogenetic lineages and over geological timescales. For example, [29] also considered another model in which speciation and extinction rates vary over time. Workers have designed birth-death models to study a variety of intrinsic and extrinsic factors that might shape diversification rates. Species age [12, 1, 32] and inherited traits [18, 25, 7, 8, 31] are two types of intrinsic factors thought to drive diversification rates, whereas environment  [5, 30] and geography [11, 21, 34] are common extrinsic factors of interest. In the end, a common goal of these models is to infer the underlying event rates given an observed phylogenetic pattern either through likelihood-based [27, 33, 24] or likelihood-free approaches [29, 41, 14, 20, 36].
Fundamentally, birth-death processes model the random arrival times of discrete events that generate or “build” a phylogenetic tree over time [29, 25]. As an alternative to this tree-based representation of the process, recent work [4] introduced an equivalent diffusion-based representation for a class of birth-death models with state-dependent rates, known as state-dependent speciation-extinction (SSE) models [25]. As noted by [4], population genetics theory has benefited immensely from diffusion-based approximations to population-based models of allele frequency change, yet diffusion-based approximations of birth-death models remain underexplored in the phylogenetics literature. Despite the widespread popularity of birth-death models among evolutionary biologists, these models recently entered a phase of intense but overdue scrutiny to better understand what the models can and cannot estimate reliably when fitted to real biological datasets [23, 28, 39, 6, 19, 22, 37, 3, 35]. This has created demand for new frameworks to understand the mathematical properties of these complex stochastic processes to guide biological research programs.
@@ 2.1 Overview of state-dependent speciation and extinction models
In general, SSE models are stochastic branching processes with state-dependent birth (speciation) and death (extinction) rates. The states can either be discrete or continuous [25, 7, 8] and can represent various things, ranging from phenotypic traits to geographical ranges [11]. Some SSE models have processes that are only defined by anagenetic process and state-dependent diversification process [25], while others have processes that are defined by both anagenetic and cladogenetic processes [11, 10] shown in Fig. 1. An anagenetic process is defined as a process of trait evolution within lineages, between branching events. In the BiSSE model [25], this corresponds to trait transition events of going from a discrete trait $A$ to another discrete trait $B$ or vice versa. These trait-dependent transition rates are encoded in the infinitesimal rate matrix $\bm{Q}$, for which the off-diagonal entry $q_{ij}$ defines the rate of transitioning from state $i$ to $j$. A cladogenetic process is defined as a process in which state transition occurs in conjunction with a branching event (with speciation) of a lineage. SSE models with anagenetic and cladogenetic events are referred to as ClaSSE models.
Figure 1: From left to right: a speciation event without cladogenetic state changes, a speciation event with cladogenetic state changes, an anagenetic state change
@@
Figure 2: An illustration of GeoSSE events on a phylogeny with range state information
@@ 2.2 Transforming a stochastic process
Note here we have used the fact that $\sigma_{ij}=0$ for $i\neq j$ to account for independent birth-death processes. The infinitesimal mean and variance parameters for $\{X(t)\}$ are given by,
@@ 2.3 Diffusion-based framework for state-dependent diversification model
In this section, we establish the framework for simulating state dynamics for state-dependent speciation and extinction models using diffusion processes. We show how to implement the framework in the ClaSSE model introduced in [10]. Then, we relate our framework to earlier research [4] using a diffusion process for the BiSSE model [25] and, later on, for the GeoSSE model [11].
@@
$\mathbb{P}_{i}^{+}\Delta t$ $=$ $S_{i}^{+}+E_{i}^{+}+Q_{i}^{+},$
$\mathbb{P}_{i}^{-}\Delta t$ $=$ $S_{i}^{-}+E_{i}^{-}+Q_{i}^{-},$
$\mathbb{P}_{i}\Delta t$ $=$ $1-\left(\mathbb{P}_{i}^{+}+\mathbb{P}_{i}^{-}\right)\Delta t,$
@@
$S_{i}^{+}$ $=$ Probability of events that lead to an increase in the number of species in state $i$ through
state-dependent speciation and speciation in conjunction with cladogenetic state change.
$E_{i}^{+}$ $=$ Probability of events that lead to an increase in the number of species in state $i$ through
extinction.
$Q_{i}^{+}$ $=$ Probability of events that lead to an increase in the number of species in state $i$ through
anagenetic state change.
$S_{i}^{-}$ $=$ Probability of events that lead to a decrease in the number of species in state $i$ through
state-dependent speciation and speciation in conjunction with cladogenetic state change.
$E_{i}^{-}$ $=$ Probability of events that lead to a decrease in the number of species in state $i$ through
extinction.
$Q_{i}^{-}$ $=$ Probability of events that lead to a decrease in the number of species in state $i$ through
anagenetic state change.
@@
To demonstrate the generality of the framework, we show the BiSSE model [25] (and similarly for the MuSSE model [8]) can be represented as a diffusion process as follows. Under the BiSSE model, species possess binary traits with values in the state space $S=\{1,2\}$. BiSSE is a special case of the ClaSSE model that, while it allows anagenetic trait transition and extinction events, its speciation events do not cause cladogenetic trait changes. That is, daughter lineages identically inherit the parent lineage state following speciation. Readers can refer to the supplementary material from [10] for its derivation. For the BiSSE model, we have
$S_{1}^{+}=\lambda_{1}N_{1}\Delta t,\>E_{1}^{+}=0,\>Q_{1}^{+}=q_{21}N_{2}\Delta t,$
$S_{1}^{-}=0,\>E_{1}^{-}=\mu_{1}N_{1}\Delta t,\>Q_{1}^{-}=q_{12}N_{1}\Delta t,$
where $\lambda_{1}$ and $\mu_{1}$ ($b_{1}$ and $d_{1}$ in [4]) are speciation and extinction rates for trait $1$, respectively. $q_{12}$ and $q_{21}$ ($\tau_{12}$ and $\tau_{21}$ in [4]) are anagenetic trait transition from $1$ to $2$ and from $2$ to $1$, respectively. Similarly, following the definitions in Eq. (LABEL:eq::generalclasse), we also have
$S_{2}^{+}=\lambda_{2}N_{2}\Delta t,\>E_{2}^{+}=0,\>Q_{2}^{+}=q_{12}N_{1}\Delta t,$
$S_{2}^{-}=0,\>E_{2}^{-}=\mu_{2}N_{2}\Delta t,\>Q_{2}^{-}=q_{21}N_{2}\Delta t,$
@@ 2.4 Diffusion-based framework for the GeoSSE model
In this section, we use the framework established in Section 2.3 for general ClaSSE models to the GeoSSE model. The procedure we apply here is also compatible with any model from the ClaSSE family. For the GeoSSE model, unlike the BiSSE model described in Section 2.3, some speciation events also cause cladogenetic state changes. Thus, following the notation used in the previous section we have,
$S_{i}^{+}$ $=$ $W_{i}^{+}+B_{i}^{+}$
$E_{i}^{+}$ $=$ $E_{i}^{+}$
$Q_{i}^{+}$ $=$ $D_{i}^{+}+E_{i}^{+}$
$S_{i}^{-}$ $=$ $W_{i}^{-}+B_{i}^{-}$
$E_{i}^{-}$ $=$ $E_{i}^{-}$
$Q_{i}^{-}$ $=$ $D_{i}^{-}+E_{i}^{-},$
where
$W_{i}^{+}$ $=$ Probability of events that lead to an increase in the number of species in range state $i$
through within-region speciation for either widespread or endemic species.
$B_{i}^{+}$ $=$ Probability of events that lead to an increase in the number of species in range state $i$
through between-region speciation for widespread species.
$E_{i}^{+}$ $=$ Probability of events that lead to an increase in the number of species in range state $i$
through extinction for either widespread species (local extinction) or endemic species
(species extinction).
$D_{i}^{+}$ $=$ Probability of events that lead to an increase in the number of species in range state $i$
through range dispersal event for endemic species.
$W_{i}^{-}$ $=$ Probability of events that lead to a decrease in the number of species in range state $i$
through within-region speciation for either widespread or endemic species.
$B_{i}^{-}$ $=$ Probability of events that lead to a decrease in the number of species in range state $i$
through between-region speciation for widespread species.
$E_{i}^{-}$ $=$ Probability of events that lead to a decrease in the number of species in range state $i$
through extinction for either widespread species (local extinction) or endemic species
(species extinction).
$D_{i}^{-}$ $=$ Probability of events that lead to a decrease in the number of species in range state $i$
through range dispersal event for endemic species.
Next, consider an $n$-region GeoSSE model where $n\in\mathbb{Z}^{+}$, we define the following state space and variable,
@@
two daughter species with ranges i and j, respectively. Note that $\text{two daughter species with ranges $i$ and $j$, respectively.Note that }b^{i}_{j}\equiv b^{j}_{i}.$
@@
Thus, both $w_{\ell}$ and $b^{i}_{j}$ determine state-dependent speciation rate, $e_{\ell}$ determines state-dependent extinction rate, and $d_{k\ell}$ and (among widespread species) $e_{\ell}$ determine the anagenetic state transition rate.
We define a stochastic process $\{N_{i}(t)\}$ with infinitesimal mean $\mu_{i}=\mathbb{E}(dN_{i}/dt)$ and variance $\sigma^{2}_{i}=var(dN_{i}/dt)$. Here, $N_{i}(t)$ represents the number of species with range state $i$ at time $t$. The infinitesimal mean $\mu_{i}$ and variance $\sigma^{2}_{i}$ follow directly from Lemma 2. We derive the transition probabilities described in Eq. (LABEL:eq:transition) in the context of the GeoSSE model, as shown in Eqs. (2.4)-(15).
@@
$\mathbb{P}_{i}^{+}\Delta t$ $=$ $W_{i}^{+}+D_{i}^{+}+B_{i}^{+}+E_{i}^{+}$
$=$ $\underbrace{\sum_{\mathclap{\begin{subarray}{c}j\in S\end{subarray}}}\sum_{\mathclap{\begin{subarray}{c}\ell\in j\\ \{\ell\}=i\end{subarray}}}N_{j}w_{\ell}\Delta t}_{W_{i}^{+}}+\underbrace{\sum_{\mathclap{\begin{subarray}{c}k\in i\end{subarray}}}\sum_{\mathclap{\begin{subarray}{c}\ell\in i\\ \ell\neq k\end{subarray}}}N_{i\backslash\{\ell\}}d_{k\ell}\Delta t}_{D_{i}^{+}}+\underbrace{\sum_{\mathclap{\begin{subarray}{c}j\in S\\ i\subset j\end{subarray}}}N_{j}b^{i}_{j\backslash i}\Delta t}_{B_{i}^{+}}+\underbrace{\sum_{\begin{subarray}{c}j\in S\\ |j\backslash i|=1\end{subarray}}\sum_{\begin{subarray}{c}\ell\in j\backslash i\end{subarray}}N_{j}e_{\ell}\Delta t}_{E_{i}^{+}}$
$\mathbb{P}_{i}^{-}\Delta t$ $=$ $W_{i}^{-}+D_{i}^{-}+B_{i}^{-}+E_{i}^{-}$ (14)
@@
$\mathbb{P}_{i}\Delta t$ $=$ $1-\left(\mathbb{P}_{i}^{+}+\mathbb{P}_{i}^{-}\right)\Delta t.$ (15)
For clarity, we provide the biogeographic interpretation on how each term in Eqs. 2.4-15 is derived and a graphical illustration of the events in Fig. 3.
@@
    $\bm{W_{i}^{+}}$. To gain a new species with range state $i$ through a within-region speciation event, the new species range $i$ must contain only region $\ell$ ($\ell\in i$ and $|i|=1$). This endemic species can undergo a speciation event with probability $w_{\ell}N_{i}$. Any species with range state $j$ that also occupies region $\ell$ can undergo a within-region speciation event with probability $w_{\ell}\sum_{j\in S}\mathbbm{1}_{i\subseteq j}N_{j}$. The total probability of this event occurring within $\Delta t$ is,
    $\sum_{\mathclap{\begin{subarray}{c}j\in S\end{subarray}}}\sum_{\mathclap{\begin{subarray}{c}\ell\in j\\ \{\ell\}=i\end{subarray}}}N_{j}w_{\ell}\Delta t.$
    As an example, in a $2-$region GeoSSE system with state space $S=\{\{A\},\{B\},\{A,B\}\}$ we have,
    $W^{+}_{\{A\}}$ $=$ $\left(N_{\{A,B\}}+N_{\{A\}}\right)w_{A}\Delta t$
    $W^{+}_{\{B\}}$ $=$ $\left(N_{\{A,B\}}+N_{\{B\}}\right)w_{B}\Delta t$
    $W^{+}_{\{A,B\}}$ $=$ $0$
@@
    $\bm{D_{i}^{+}}$. To gain a new species with range state $i$ through a dispersal event, the species adds the new region $\ell$ to its ancestral range. Species are always widespread immediately following dispersal. The total probability of this event occurring within $\Delta t$ is,
    $\sum_{\mathclap{\begin{subarray}{c}k\in i\end{subarray}}}\sum_{\mathclap{\begin{subarray}{c}\ell\in i\\ \ell\neq k\end{subarray}}}N_{i\backslash\{\ell\}}d_{k\ell}\Delta t.$
    As an example, in a $2-$region GeoSSE system with state space $S=\{\{A\},\{B\},\{A,B\}\}$ we have,
    $D^{+}_{\{A\}}$ $=$ 00 0
    $D^{+}_{\{B\}}$ $=$ 00 0
    $D^{+}_{\{A,B\}}$ $=$ $N_{\{B\}}d_{BA}\Delta t+N_{\{A\}}d_{AB}\Delta t.$
@@
    $\bm{B_{i}^{+}}$. To gain a new species with range state $i$ through a between-region speciation event, the new species can be either endemic or widespread $|i|>0$ that originated from a widespread ancestral species with larger range state $j$ ($i\subset j$). In general, we have no information of whether the new species occurs in left or right lineage following a speciation event, so we do not consider the orientation. The total probability of this event occurring within $\Delta t$ is,
    $\sum_{\mathclap{\begin{subarray}{c}j\in S\\ i\subset j\end{subarray}}}N_{j}b^{i}_{j\backslash i}\Delta t.$
    As an example, in a $2-$region GeoSSE system with state space $S=\{\{A\},\{B\},\{A,B\}\}$ we have,
    $B^{+}_{\{A\}}$ $=$ $N_{\{A,B\}}b^{A}_{B}\Delta t$
    $B^{+}_{\{B\}}$ $=$ $N_{\{A,B\}}b^{B}_{A}\Delta t$
    $B^{+}_{\{A,B\}}$ $=$ $0$
@@
    $\bm{E_{i}^{+}}$. To gain a new species with range state $i$ through a local extinction event, the ancestral species must have a larger range state $j$ with size that differs by 1 from the new species’ range state $i$ such that $|j\backslash i|=1$. The total probability of this event occurring within $\Delta t$ is,
    $\sum_{\begin{subarray}{c}j\in S\\ |j\backslash i|=1\end{subarray}}\sum_{\begin{subarray}{c}\ell\in j\backslash i\end{subarray}}N_{j}e_{\ell}\Delta t.$
    As an example, in a $2-$region GeoSSE system with state space $S=\{\{A\},\{B\},\{A,B\}\}$ we have,
    $E^{+}_{\{A\}}$ $=$ $N_{\{A,B\}}e_{B}\Delta t$
    $E^{+}_{\{B\}}$ $=$ $N_{\{A,B\}}e_{A}\Delta t$
    $E^{+}_{\{A,B\}}$ $=$ $0$
@@
    $\bm{D_{i}^{-}}$. To lose a species with range state $i$ through a dispersal event, the species must disperse to a new region. The species count remains unchanged if the species already occupies all regions ($|i|=|R|$). The total probability of this event occurring within $\Delta t$ is,
@@
    As an example, in a $2-$region GeoSSE system with state space $S=\{\{A\},\{B\},\{A,B\}\}$ we have,
    $D^{-}_{\{A\}}$ $=$ $N_{\{A\}}d_{AB}\Delta t$
    $D^{-}_{\{B\}}$ $=$ $N_{\{B\}}d_{BA}\Delta t$
    $D^{-}_{\{A,B\}}$ $=$ $0$
@@
    $\bm{B_{i}^{-}}$. To lose a species with range state $i$ through a between-region speciation event, the species must be widespread and undergo a speciation event that gives rise to a new species in state $j$ with smaller range state size ($|j|<|i|$). The factor of 1/2 corrects for double-counting the new species with range $j$ being either the left daughter or right daughter lineage. The total probability of this event occurring within $\Delta t$ is,
    $\sum_{\begin{subarray}{c}j\in S\\ |j|<|i|\end{subarray}}\frac{1}{2}N_{i}b^{j}_{i\backslash j}\Delta t.$
    As an example, in a $2-$region GeoSSE system with state space $S=\{\{A\},\{B\},\{A,B\}\}$ we have,
    $B^{-}_{\{A\}}$ $=$ 00 0
    $B^{-}_{\{B\}}$ $=$ 00 0
    $B^{-}_{\{A,B\}}$ $=$ $\frac{1}{2}N_{\{A,B\}}\left(b^{A}_{B}+b^{B}_{A}\right)\Delta t.$
@@
    $\bm{E_{i}^{-}}$. To lose a species with range state $i$ through a local extinction event, a species must undergo an extinction event in one of its regions. If the species is endemic, this event leads to total extinction of the species. The total probability of this event occurring within $\Delta t$ is,
    $\sum_{\begin{subarray}{c}\ell\in R\\ \ell\in i\end{subarray}}N_{i}e_{\ell}\Delta t.$
    As an example, in a $2-$region GeoSSE system with state space $S=\{\{A\},\{B\},\{A,B\}\}$ we have,
    $E^{-}_{\{A\}}$ $=$ $N_{\{A\}}e_{A}\Delta t$
    $E^{-}_{\{B\}}$ $=$ $N_{\{B\}}e_{B}\Delta t$
    $E^{-}_{\{A,B\}}$ $=$ $N_{\{A,B\}}\left(e_{A}+e_{B}\right)\Delta t.$
The next section uses Eqs. 2.4-15 to define the stochastic process $\{\Pi_{i}(t):t>0\}$ that models the frequency of species in range state $i$ at time t. The infinitesimal mean $\mu_{\Pi_{i}}$ and variance $\sigma^{2}_{\Pi_{i}}$ follow directly from Lemma 3.
Figure 3: Graphical illustrations of probabilities of events following Eq. (2.4) shown in (a), and Eq. (14) shown in (b) for a 2-region GeoSSE system with state space $S=\{\{A\},\{B\},\{A,B\}\}$. $N_{i}$ represents the number of species with range state $i\in S$. An incoming arrow into $N_{i}$ compartment means there is an increase in species count with range state $i$ and an outgoing arrow from $N_{i}$ means there is a decrease in species count with range state $i$. All the events and arrows are color-coded accordingly.